OpenStudy (anonymous):

Find a Cartesian equation for the curve and identify it. r = 6 tan θ sec θ

OpenStudy (anonymous):

\[\begin{align*} r&=6\tan\theta\sec\theta\\\\ r&=\frac{6\dfrac{\sin\theta}{\cos\theta}}{\cos\theta}\\\\ r\cos\theta&=6\frac{\sin\theta}{\cos\theta}\\\\ r\cos\theta&=6\frac{\sin\theta}{\cos\theta}\cdot\frac{r}{r}\\\\ x&=6\frac{y}{x}\\\\ y&=\frac{1}{6}x^2 \end{align*}\]